# via Irnerio 46, I-40126 Bologna, Italy

### v2

@@ 2.1 Notation
where $R=\sqrt{s_{12}^{2}+(m_{4}^{2}-m_{3}^{2})^{2}-2s_{12}(m_{4}^{2}+m_{3}^{2})}$. As in Henn:2014lfa , the square root $R$ can be rationalized by transforming into the variables $(x,y,z)$ defined by
@@ 3.1 Canonical form of the differential equations
and 15 in the rationalized variables, since we set $m_{3}^{2}=1$
@@ 3.2 Construction of pure bases
If this system is solved, the new candidates defined as $\vec{I}^{\text{\tiny MC}}=\tilde{T}^{\text{\tiny MC}}\vec{G}^{\text{\tiny MC}}$ satisfy canonical DEs in the maximal cut [^6] The notation MC (LS) here and later on stands for symbolizing quantities in the maximal cut (lower sectors).
@@
In the former, the constant leading singularities of one-loop and two-loop pure candidates are used as building blocks, which combined create the leading singularity of a three-loop candidate, in a graphical and heuristic approach. This method has been shown to work well with massless propagators, and in our study it helped us to identify pure candidates including also the necessary lower-sector contributions. For the latter method, we used two approaches. Firstly, we used the DLogBasis package Henn:2020lye , which brings the 4-dimensional integrand of a Feynman integral into $d\log$ form using the spinor helicity formalism. In order to employ the latter for the massive momenta, we used a two-variable simplified differential equations parametrization Papadopoulos:2014lla to re-express them in terms of massless momenta, as proposed in Canko:2021hvh . Secondly, we exploited the loop-by-loop Baikov representation of Feynman integrals Frellesvig:2017aai for some sectors, trying to put the resulting d-dimensional integrand into a dlog-form for the Baikov variables, first in the maximal cut and later by investigating sub-maximal contributions.
@@ 4.1 Analytic results in terms of multiple polylogarithms
and the Mathematica package DiffExp Hidding:2020ytt for the evaluation of the GPLs at the same point with the same precision. After inserting these numerical values, the coefficient of $\varepsilon^{w}$ in Eq. (4.1) is composed of a linear combination of complex numbers with undetermined rational coefficients (the boundary constant $\vec{b}_{w}$) and a complex number obtained by summing all other terms at this weight (including $\vec{I}^{(w)}$). In this way, the problem of determining the unknown boundary conditions of the pure MIs is reduced to that of finding integer coefficients for a linear combination of real numbers [^7] Formally, we multiplied the equation by the minimum common multiple of the rational coefficients and treated separately its real and the imaginary part, resulting in two equations.. We solved this problem employing the PSLQ algorithm pslq .
@@ 4.2 Numerical results using DiffExp
$\mathrm{PL}_{1}$ 126 $s$ 136 s<

### v3

@@ 2.1 Notation
where
$R=\sqrt{s_{12}^{2}+(m_{4}^{2}-m_{3}^{2})^{2}-2s_{12}(m_{4}^{2}+m_{3}^{2})}.$ (3)
As in Henn:2014lfa , the square root $R$ can be rationalized by transforming into the variables $(x,y,z)$ defined by
@@ 3.1 Canonical form of the differential equations
where $R$ is the square root defined in Eq. (3), and 15 in the rationalized variables, since we set $m_{3}^{2}=1$
@@ 3.2 Construction of pure bases
The differential equation is in general difficult to solve. In practice, one determines each entry separately, by observing that the diagonal ones correspond to the normalisation of the MIs and the off-diagonal ones to the contribution of the MIs to the derivatives of the others. Therefore, one has to build linear combinations of appropriately normalised master integrals. The new candidates defined as $\vec{I}^{\text{\tiny MC}}=\tilde{T}^{\text{\tiny MC}}\vec{G}^{\text{\tiny MC}}$ satisfy canonical DEs in the maximal cut [^6] The notation MC (LS) here and later on stands for symbolizing quantities in the maximal cut (lower sectors).
@@
In the former, the constant leading singularities of one-loop and two-loop pure candidates are used as building blocks, which combined create the leading singularity of a three-loop candidate, in a graphical and heuristic approach. This method has been shown to work well with massless propagators, and in our study it helped us to identify pure candidates including also the necessary lower-sector contributions. For the latter method, we used two approaches. Firstly, we used the DLogBasis package Henn:2020lye , which brings the 4-dimensional integrand of a Feynman integral into $d\log$ form using the spinor helicity formalism. In order to employ the latter for the massive momenta, we used a two-variable simplified differential equations parametrization Papadopoulos:2014lla to re-express them in terms of massless momenta, as proposed in Canko:2021hvh . Secondly, we exploited the loop-by-loop Baikov representation of Feynman integrals Baikov:1996iu ; Baikov:1996rk for some sectors, using the package Baikov.m Frellesvig:2017aai , trying to put the resulting d-dimensional integrand into a dlog-form for the Baikov variables, first in the maximal cut and later by investigating sub-maximal contributions.
@@ 4.1 Analytic results in terms of multiple polylogarithms
and the Mathematica package DiffExp Hidding:2020ytt for the evaluation of the GPLs at the same point with the same precision. After inserting these numerical values, the coefficient of $\varepsilon^{w}$ in Eq. (4.1) is composed of a linear combination of complex numbers with undetermined rational coefficients (the boundary constant $\vec{b}_{w}$) and a complex number obtained by summing all other terms at this weight (including $\vec{I}^{(w)}$). In this way, the problem of determining the unknown boundary conditions of the pure MIs is reduced to that of finding integer coefficients for a linear combination of real numbers [^7] Formally, we multiplied the equation by the minimum common multiple of the rational coefficients and treated separately its real and the imaginary part, resulting in two equations.. We solved this problem employing the PSLQ algorithm pslq , as implemented in Mathematica.
@@ 4.2 Numerical results using DiffExp
$\mathrm{PL}_{1}$ 126 $s$ 136 $s$
$\mathrm{PT}_{4}$ 226 $s$ 246 $s$
Table 4: Average time needed by DiffExp to integrate a segment with 40 digits precision, per family, and per set of kinematic variables used for the DEs.
We conclude this section by commenting on the performance of DiffExp for the two sets of kinematic invariants. In the table 3 we show the average time needed to compute a single point for the two sets of kinematic invariants. As can be seen from the results, evaluating the integrals in the variables $(x,y,z)$ is considerably faster. However, the time per kinematic point is not the only criterion on which the performance of DiffExp should be judged. Since DiffExp internally divides the path from the base-point to the target point into segments, which are then integrated one after the other, one should also look at the time needed to integrate a single segment. We computed the average time per segment for all the families. The results, shown in table 4, paint the opposite picture compared to table 3. This is due to the fact that the DEs in the variables $(s_{12},s_{23},m_{3}^{2},m_{4}^{2})$ are less complicated than those in $(x,y,z)$, as can be deduced from the size of the files in the ancillaries. Therefore, the integration of the differential equation is faster. However, the phase-space has a simpler topology when the square root is rationalized, hence DiffExp splits the integration path into fewer sub-segments. Finally, with an eye towards phenomenology, we remark that constructing a grid in a smart way one could integrate each point using only one segment, in which case the timings in table 4 would approximately be the time needed to compute one point.
5 Conclusions
In this paper, we undertook a study of a ladder-box, a tennis-court, and two reducible ladder-box-like three-loop integral families with massless propagators and four external particles, where two of the external particles are massive but bear different masses. We constructed pure bases and derived canonical DEs, which we subsequently solved on physical kinematics up to order six in the dimensional regulator. This was done analytically in terms of MPLs and numerically utilizing DiffExp, verifying the validity of our results by comparisons with AMFlow, finding perfect agreement in all cases. These families are of significant phenomenological interest, as they are relevant for the study of the production of two massive vector bosons at N3LO precision in QCD.
We expect that the analytic solution can still be optimized, for example by identifying additional relations between the MPLs that appear in it. For this reason, future work will involve the complete decomposition into Lyndon words and the application of symbol algebra to reveal relations between the existing MPLs, so that solutions can be expressed in terms of a minimal set of independent MPLs. Moreover, we plan to implement the evaluation of the analytic solution using more conventional tools, such as a C++ (FORTRAN) code suitable for fast and parallel numerical evaluations through GiNaC ginac (handyG Naterop:2019xaf ). Alternatively, the efficiency of the numerical solution of the DiffExp solution can be enhanced by creating an optimized grid of numerical solutions in multiple phase-space points, as demonstrated in Abreu:2020jxa . Whether pursuing one or the other path, we expect that our solutions will prove efficient for the future computation of scattering amplitudes involving these integrals.
The study of the remaining planar (and non-planar) three-loop four-point integral families with two off-shell legs represents a logical subsequent step in the same direction of this work. Based on the existing studies of analogous problems Canko:2021hvh ; Long:2024bmi , it is reasonable to expect that the remaining planar families can be expressed in terms of MPLs too. Consequently, a comparable approach to the one employed in this study can be effectively applied to address these families too. Moreover, once all the planar families have been computed, it would be worthwhile to investigate the adjacency conditions between the letters of their alphabets and the possibility of finding a cluster algebra describing them, as has already been done for the corresponding planar families with one off-shell leg Chicherin:2020umh ; Aliaj:2024zgp .
Acknowledgements.
We would like to thank Matteo Becchetti, Vsevolod Chestnov, Tiziano Peraro, and Simone Zoia for useful discussions and comments on the draft. This work was supported by the European Research Council (ERC) under the European Union’s Horizon Europe research and innovation program grant agreement 101040760, High-precision multi-leg Higgs and top physics with finite fields (ERC Starting Grant FFHiggsTop).
Appendix A Description of the ancillary files
The ancillary files can be downloaded from canko_2024_14284044 . The notation in the files translates to that of the article as follows:
sij = $s_{ij}$, FFG[Fam,{a1,...,a15}] = $G^{Fam}_{a_{1},\dots,a_{15}}$,
mi2 = $m_{i}^{2}$, W[i],W[r[1],i] = $w_{i}$ (respectively $\bar{w}_{i}$),
eps = $\epsilon$, dlog[W[i]] = $\mathrm{d}\log(w_{i})$ (respectively $\mathrm{d}\log(\bar{w}_{i})$),
r[1] = $R$, G[a1,…,an,x] = $\mathcal{G}(a_{1},\dots,a_{n};X)$,
Pi = $\pi$, Zeta[3] = $\zeta_{3}$,
Zeta[5] = $\zeta_{5}$, I = $\mathrm{i}$.
The folder GlobalDefinitions contains the definitions of the alphabet for both choices of variables, in the format
[…]
We remark that in general $w_{i}\neq\bar{w}_{i}$, even though both are called W[i] in the ancillary files.
For each family, we provide two folders:
  • MIs_DEs.
  • AnalyticSolution.
The first one contains five files:
  • masters.m: the basis of pure candidates, in terms of FFG[Fam,{a1,...,a15}], eps and external invariants (s12, s23, m32, m42).
  • boundaries.m and dAtilde.m: the values of the pure candidates at the boundary point of Eq. (27) and the matrix $\tilde{A}$ from Eq. (12), both in the Mandelstam variables $\left(s_{12},s_{23},m_{3}^{2},m_{4}^{2}\right)$.
  • boundariesxyz.m and dAtildexyz.m: the same as the previous two files, but in the rationalized variables $(x,y,z)$.
The second folder contains the zip file with the analytic solution, in terms of the MPLs G[a1,…,an,x], the transcendental constants Pi, Zeta[3], Zeta[5] and the dimensional regulator eps.
Lastly, we provide the codes DiffExpRun.wl, which computes numerically the pure basis elements using DiffExp. In order to use it, one should first modify the variable PathToDiffExp in the file DiffExpRun.wl appropriately, with the path to the DiffExp.m file on their machine. Then the only input needed is a file with the physical point where one wants to evaluate the integrals, in the same format as testpoint.m (which is provided in the ancillaries). The command to run the computation is then
math -script DiffExpRun.wl -family ’fam’ -vars ’vars’/
-filein "point.m"/
[…]
where ’fam’ is the integral family one wants to compute, ’vars’ is either ’sij’ or ’xyz’ and the results are stored as a list in output.m. On the first use, the file will compute the partial derivatives of the pure basis from the matrix $\tilde{A}$ and store them in the fam/MIs_DEs folder.
Appendix B Genuinely new sectors
In this appendix we collect the pure candidates for the genuinely new sectors, referring to their position on the basis of the corresponding integral family (see ancillary files canko_2024_14284044 ). In particular, concerning the sector definition, we are using the notation
$\text{SFam}[a_{1},\ldots,a_{15}],$
[…]
$I_{\text{PS}}^{\text{Fam}},$
[…]
In total, 50 new sectors were identified, consisting of 153 MIs. In the following we present a classification of the sectors based on the number of propagators they contain, commencing with those having five propagators and concluding with those comprising ten.
Five-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[0,0,1,0,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{133}^{\text{PL1}},\,I_{134}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,0,1,0,0,0,0,0,0,1,1,0,1]
$\{I_{126}^{\text{PL1}},\,I_{127}^{\text{PL1}},\,I_{128}^{\text{PL1}}\}$
Six-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[0,0,1,0,1,0,0,0,1,0,0,1,1,0,1]
$\{I_{122}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,0,0,0,0,0,0,0,1,1,1,0,1]
$\{I_{102}^{\text{PL1}},\,I_{103}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,0,1,0,0,0,0,0,1,1,1,0,1]
$\{I_{87}^{\text{PL1}},\,I_{88}^{\text{PL1}},\,I_{89}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,0,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{106}^{\text{PL1}},\,I_{107}^{\text{PL1}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,0,0,0,0,0,1,0,0,1,1,0,1]
$\begin{split}\{I_{4}^{\text{RL2}},\,I_{5}^{\text{RL2}},\,I_{6}^{\text{RL2}},\,I_{7}^{\text{RL2}},\,I_{8}^{\text{RL2}},\,I_{9}^{\text{RL2}}\}\end{split}$
Sector $\mathbf{F_{123}}$[0,1,1,0,1,0,0,0,0,0,0,1,1,0,1]
$\{I_{104}^{\text{PL1}},\,I_{105}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,0,0,0,0,1,1,0,1,1,0]
$\{I_{136}^{\text{PT4}},\,I_{137}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{77}^{\text{PL1}},\,I_{78}^{\text{PL1}},\,I_{79}^{\text{PL1}},\,I_{80}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,0,0,0,0,1,0,0,0,1,1]
$\{I_{123}^{\text{PT4}},\,I_{124}^{\text{PT4}},\,I_{125}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,0,0,0,0,1,1,0,1]
$\{I_{110}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,0,0,0,0,1,0,1,1,1,0,1]
$\{I_{108}^{\text{PL1}},\,I_{109}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,0,0,0,0,1,0,0,0,1,0,1]
$\{I_{13}^{\text{RL1}}\}$
Sector $\mathbf{F_{132}}$[0,1,1,1,0,0,0,0,1,0,0,0,1,0,1]
$\{I_{14}^{\text{RL2}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,1,0,1,0,0,0,0,1,1,0,1]
$\{I_{81}^{\text{PL1}},\,I_{82}^{\text{PL1}},I_{83}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,0,0,0,0,1,1,0,0,1,1]
$\{I_{155}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[1,0,0,1,0,1,0,0,0,0,1,0,1,1,0]
$\{I_{114}^{\text{PT4}},I_{115}^{\text{PT4}},I_{116}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,0,0,0,0,0,1,0,1,1,1]
$\{I_{104}^{\text{PT4}},I_{105}^{\text{PT4}},I_{106}^{\text{PT4}},I_{107}^{\text{PT4}}\}$
Seven-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,0,1,0,0,0,0,0,1,0,1,1,1,0,1]
$\{I_{71}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,0,0,0,1,1,1,0,1]
$\{I_{66}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,1,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{4}^{\text{RL1}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,0,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{2}^{\text{RL2}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,1,0,0,0,0,1,1,0,1]
$\{I_{64}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,1,0,1,0,0,0,0,1,0,1,0,1]
$\{I_{81}^{\text{PT4}},\,I_{82}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,1,0,0,1,0,1,0,0,0,1,1,1,0,1]
$\{I_{61}^{\text{PL1}},\,I_{62}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,0,0,0,0,1,1,0,0,1,1]
$\{I_{95}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,1,0,0,0,1,0,0,1,1,0]
$\{I_{87}^{\text{PT4}},\,I_{88}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,1,0,0,0,0,1,0,1,0,1]
$\{I_{74}^{\text{PT4}},\,I_{75}^{\text{PT4}},\,I_{76}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{52}^{\text{PL1}},\,I_{53}^{\text{PL1}},\,I_{54}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,1,1,0,1,0,0,0,1,0,0,1,1,0,1]
$\{I_{57}^{\text{PL1}},\,I_{58}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,1,0,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{37}^{\text{PL1}},\,I_{38}^{\text{PL1}},\,I_{39}^{\text{PL1}},\,I_{40}^{\text{PL1}},\,I_{41}^{\text{PL1}},\,I_{42}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,1,0,0,0,1,1,0,1,1,0]
$\{I_{91}^{\text{PT4}},\,I_{92}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,0,0,0,0,1,1,0,1,1,1]
$\{I_{67}^{\text{PT4}},\,I_{68}^{\text{PT4}},\,I_{69}^{\text{PT4}},\,I_{70}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,0,0,0,0,1,0,0,1,1,1]
$\begin{split}\{I_{58}^{\text{PT4}},\,I_{59}^{\text{PT4}},\,I_{60}^{\text{PT4}},\,I_{61}^{\text{PT4}},\,I_{62}^{\text{PT4}}\}\end{split}$
Sector $\mathbf{F_{123}}$[1,0,1,1,0,1,0,0,0,1,0,0,1,1,0]
$\begin{split}\{&I_{77}^{\text{PT4}},\,I_{78}^{\text{PT4}}\}\end{split}$
Eight-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[0,1,0,0,1,0,1,0,1,0,1,1,1,0,1]
$\{I_{34}^{\text{PL1}},\,I_{35}^{\text{PL1}},\,I_{36}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,1,0,1,0,0,0,0,1,1,0,1]
$\{I_{22}^{\text{PL1}},\,I_{23}^{\text{PL1}},\,I_{24}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,1,1,0,0,0,0,1,1,0,0,1,1]
$\{I_{36}^{\text{PT4}},\,I_{37}^{\text{PT4}},\,I_{38}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,0,0,1,0,1,1,1,0,1]
$\{I_{25}^{\text{PL1}},\,I_{26}^{\text{PL1}},\,I_{27}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,1,0,0,0,0,1,0,1,0,1,1,1,0,1]
$\{I_{28}^{\text{PL1}},\,I_{29}^{\text{PL1}},\,I_{30}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,0,0,0,1,0,1,0,0,1,1,0,1]
$\{I_{17}^{\text{PL1}},\,I_{18}^{\text{PL1}},\,I_{19}^{\text{PL1}},\,I_{20}^{\text{PL1}},\,I_{21}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,0,1,1,1,0,0,0,1,1,0,1,1,1]
$\{I_{39}^{\text{PT4}},\,I_{40}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,0,0,0,0,1,1,0,1,1,1]
$\{I_{21}^{\text{PT4}},\,I_{22}^{\text{PT4}},\,I_{23}^{\text{PT4}},\,I_{24}^{\text{PT4}},\,I_{25}^{\text{PT4}},\,I_{26}^{\text{PT4}}\}$
Sector $\mathbf{F_{123}}$[1,1,1,1,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{1}^{\text{RL1}}\}$
Sector $\mathbf{F_{132}}$[1,1,1,1,0,0,0,0,1,0,0,1,1,0,1]
$\{I_{1}^{\text{RL2}}\}$
Nine-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,1,1,0,1,0,1,0,1,0,0,1,1,0,1]
$\{I_{6}^{\text{PL1}},\,I_{7}^{\text{PL1}},\,I_{8}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[0,0,1,1,1,1,0,0,0,1,1,0,1,1,1]
$\{I_{10}^{\text{PT4}},\,I_{11}^{\text{PT4}},\,I_{12}^{\text{PT4}},\,I_{13}^{\text{PT4}}\}$
Ten-Propagator Pure Candidates
Sector $\mathbf{F_{123}}$[1,1,1,0,1,0,1,0,1,0,1,1,1,0,1]
$\{I_{1}^{\text{PL1}},\,I_{2}^{\text{PL1}},\,I_{3}^{\text{PL1}},\,I_{4}^{\text{PL1}},\,I_{5}^{\text{PL1}}\}$
Sector $\mathbf{F_{123}}$[1,0,1,1,1,1,0,0,0,1,1,0,1,1,1]
$\{I_{1}^{\text{PT4}},\,I_{2}^{\text{PT4}},\,I_{3}^{\text{PT4}},\,I_{4}^{\text{PT4}},\,I_{5}^{\text{PT4}}\}$